interconverting compound SI units

Interconverting compound Sl units Convert the following measurement. -0 -4 g mg 4.9 x 10 L dL X Exnlanation 1I
Answer #1

Conversion to be used:

1 g = 1000 mg

1 L = 10 dL

we have:

4.9*10^-4 g/L

= 4.9*10^-4 * 1000 mg / 10 dL

= 4.9*10^-4 * 100 mg / dL

= 4.9*10^-2 mg / dL

Answer: 4.9*10^-2